OpenStudy (anonymous):

find the slope intercept form of the equation y-4=-5/6(x-6)

OpenStudy (anonymous):

\[ y-4=\frac{-5}{6}(x-6) \rightarrow y-4=\frac{-5}{6}x+\frac{-5}{6} \times 6 \] \[ y-4=\frac{-5}{6}x+(-5) \rightarrow y=\frac{-5}{6}x-5+4 \rightarrow y=-\frac{5}{6}x-1\] Thus \[y=-\frac{5}{6}x-1\] is in the required slope -intercept form Where slope m=-5/6 and y-intercept =c=-1 @sierrajohnson95